# Fusing Global and Local: Transformer-CNN Synergy for Next-Gen Current Estimation

### v1

@@ Abstract
This paper presents a hybrid model combining Transformer and CNN for predicting the current waveform in signal lines. Unlike traditional approaches such as current source models, driver linear representations, waveform functional fitting, or equivalent load capacitance methods, our model does not rely on fixed simplified models of standard-cell drivers or RC loads. Instead, it replaces the complex Newton iteration process used in traditional SPICE simulations, leveraging the powerful sequence modeling capabilities of the Transformer framework to directly predict current responses without iterative solving steps. The hybrid architecture effectively integrates the global feature-capturing ability of Transformers with the local feature extraction advantages of CNNs, significantly improving the accuracy of current waveform predictions.
Experimental results demonstrate that, compared to traditional SPICE simulations, the proposed algorithm achieves an error of only 0.0098. These results highlight the algorithm’s superior capabilities in predicting signal line current waveforms, timing analysis, and power evaluation, making it suitable for a wide range of technology nodes, from 40nm to 3nm.
@@
— Signal line, RC network, current response, simulation, Transformer, CNN , Deep learning
@@ I Introduction
Despite significant efforts to improve the efficiency and accuracy of predicting signal line RC responses, traditional driver-side modeling approaches often struggle to cope with the complexities of high-order RC loads. To simplify these intricate RC networks, researchers typically employ model order reduction (MOR) techniques, reducing them to effective capacitances or $\pi$-models. [29][30]However, these methods rely on the linear characteristics of RC circuits, making it easy to overlook subtle variations and complex coupling effects within the signal lines. Consequently, achieving high precision while enhancing computational efficiency remains a critical research challenge in the field of RC response prediction.
@@
This paper presents a hybrid model combining Transformer and CNN for predicting the current waveform in signal lines. Unlike traditional approaches such as current source models, driver linear representations, waveform functional fitting, or equivalent load capacitance methods, our model does not rely on fixed simplified models of standard-cell drivers or RC loads. Instead, it replaces the complex Newton iteration process used in traditional SPICE simulations, leveraging the powerful sequence modeling capabilities of the Transformer framework to directly predict current responses without iterative solving steps. The hybrid architecture effectively integrates the global feature-capturing ability of Transformers with the local feature extraction advantages of CNNs, significantly improving the accuracy of current waveform predictions.
@@
II Data Acquisition and Preprocessing
II-A Data Source and Simulation Configuration
II-A1 Simulation Environment Setup
@@ II-A2 Stimulation Configuration
II-B Data Preprocessing Methods
II-B1 Data Sources and Feature Definition
@@
  • RC Parameters: The extracted parasitic resistance ($R$) and capacitance ($C$) from the SPEF netlist form a two-dimensional matrix $M\in\mathbb{R}^{N\times 2}$, where $N$ represents the number of circuit nodes.
@@ II-B2 Feature Normalization
  • RC Parameters: Z-score normalization is applied to both resistance ($R$) and capacitance ($C$) values to standardize them with a mean of 0 and a standard deviation of 1:
    $R^{\prime}_{i}=\frac{R_{i}-\mu_{R}}{\sigma_{R}},\quad C^{\prime}_{j}=\frac{C_{j}-\mu_{C}}{\sigma_{C}}$ (2)
    In these equations, $R_{i}$ and $C_{j}$ are the original resistance and capacitance values, while $\mu_{R}$ and $\mu_{C}$ denote their respective means, and $\sigma_{R}$ and $\sigma_{C}$ their standard deviations. Standardizing RC parameters is critical to ensure that features with different scales contribute equally to the model training, thereby reducing numerical instability.
@@
II-C Detailed Procedure for SPEF Netlist Parsing and RC Parameter Flattening
This section provides a detailed explanation of the RC parameter extraction and flattening process based on SPEF netlists, covering data preprocessing, the extraction of key information, and the final mathematical representation.
II-C1 Data Preprocessing and Key Information Definition
According to SPEF syntax, four categories of key information are extracted for each RC element. The details are summarized in the following table:
During this process, the extracted information is appropriately normalized. For example, node positions are converted to normalized coordinates, and electrical values undergo Z-score normalization. These steps help eliminate differences in measurement scales and prepare the data for subsequent feature fusion.
II-D RC Network Modeling and Feature Representation
To precisely characterize the electrical behavior of signal interconnects, we propose a graph-based fine-grained RC network modeling method. The core elements are as follows:
1. Network Topology Definition
Each signal wire is modeled as a weighted graph $\mathcal{G}_{i}=(\mathcal{V}_{i},\mathcal{E}_{i})$, where:
  • The vertex set $\mathcal{V}_{i}=\{n_{i1},n_{i2},...,n_{iN}\}$ represents spatially discretized nodes along the wire.
  • The edge set $\mathcal{E}_{i}$ consists of two types of elements:
    • Resistor edges $R_{jk}\in\mathcal{E}_{R}$: Parasitic resistance between adjacent nodes.
    • Capacitor edges $C_{jk}\in\mathcal{E}_{C}$: Ground or coupling capacitance of nodes.
2. Node Capacitance Modeling
The total ground capacitance of each node $n_{ij}$ is computed as follows:
$C_{ij}=\underbrace{\sum_{k=1}^{M_{p}}c^{\text{para}}_{ijk}}_{\text{Parasitic Capacitance}}+\underbrace{\sum_{l=1}^{M_{e}}c^{\text{exp}}_{ijl}}_{\text{Explicit Capacitance}}$
where $M_{p}$ is the number of parasitic capacitances extracted from process data, and $M_{e}$ is the number of explicit capacitors defined in the netlist.
3. Effective Resistance Modeling
Considering the impact of connection topology, a degree-weighted distribution strategy is applied:
$R_{ij}=\sum_{k\in\mathcal{N}(j)}\frac{R_{jk}}{d_{j}+\epsilon}$
@@ 3. Effective Resistance Modeling
  • $\mathcal{N}(j)$: The set of neighboring nodes of $n_{ij}$.
  • $d_{j}$: The degree (number of connected resistors) of node $j$.
  • $\epsilon$: A small smoothing factor (default $10^{-6}$) to avoid division by zero.
4. Node Feature Encoding
To enable batch processing by graph neural networks, each node is encoded as a 4-dimensional feature vector:
$\mathbf{x}_{ij}=\left[\phi(i),\psi(j),\frac{C_{ij}}{C_{\text{norm}}},\frac{R_{ij}}{R_{\text{norm}}}\right]\in\mathbb{R}^{4}$
@@ 4. Node Feature Encoding
  • $\phi(\cdot)$: Network-level identifier extracted from the netlist, indicating which signal net the node belongs to.
  • $\psi(\cdot)$: Node-level identifier extracted from the netlist, representing the node’s relative position or index within the signal net.
  • $C_{\text{norm}},R_{\text{norm}}$: Normalization constants derived from training data statistics.
Technical Advantages
1. Physical Interpretability: Preserves original SPEF netlist topology and device parameters. 2. Computational Efficiency: Degree-weighted strategy reduces $O(N^{2})$ resistance mapping to $O(N)$ feature computation.
This encoding scheme effectively captures the essential physical and electrical properties of each RC element, transforming them into a unified feature vector. Such a representation is crucial for facilitating robust multi-modal feature fusion and ensuring high-quality inputs for subsequent modeling tasks.
III Hybrid Model Architecture Design
III-A Overall Architecture Overview
@@
III-B Input Embedding Layer
@@
III-C CNN Architecture Design
III-C1 Input Feature Adaptation
@@
III-C2 Convolutional Feature Extraction
@@
III-C3 Feature Space Mapping
@@
III-D Encoder Architecture and Feature Fusion Strategy
@@
III-D1 Input Feature Preprocessing
@@
III-D2 Multi-Head Self-Attention Mechanism
@@
III-D3 Residual Connection and Layer Normalization
@@
III-D4 Feedforward Neural Network
@@
III-D5 Second Residual Connection and Normalization
@@ III-D6 Encoder Output
III-D7 Feature Fusion via Dot Product
@@
III-D8 Decoder Input Projection
@@
III-E Transformer Decoder Architecture
@@ III-E1 Input Definition
III-E2 Masked Multi-Head Self-Attention
@@
III-E3 Encoder-Decoder Cross-Attention
@@
III-E4 Feedforward Neural Network
@@ III-E5 Summary
IV Results
IV-A Scalability Across Network Sizes
@@
IV-B Accuracy of Voltage Response Prediction
@@
IV-C Training time and Infer time Comparison
@@
V Future Challenges and Directions for our model
VI Future Challenges and Directions
@@
VI-A Limited Generalization Across Topologies
@@
VI-B Quadratic Complexity of Transformer Inference
@@
VI-C Temporal Alignment and Stability Under Noise
@@
VI-D Lack of Physical Interpretability and Control

### v2

@@ Abstract
This paper proposes an end-to-end method for fast voltage waveform prediction based on a hybrid CNN–Transformer architecture. Unlike traditional approaches that iteratively solve the complete RC topology, the proposed method innovatively incorporates the transfer function as a high-level input feature, reducing modeling complexity from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n)$ and significantly lowering computational overhead.
To enable generalization to arbitrary-order transfer functions, we exploit the physical property that any high-order transfer function can be decomposed into a sum of first-order and repeated-pole components. Accordingly, we design a two-stage sub-model: a base model predicts the response of the dominant first-order term, while a compensation module consisting of a single correction network directly outputs point-wise correction values. These corrections are additively applied to the base response to generate the complete high-order prediction.
The proposed method eliminates the need for iterative computation or manual simplification, achieving both high modeling efficiency and generalization capability. Experimental results show that, compared to HSPICE, our approach achieves a root mean square error (RMSE) as low as 0.0098 and maintains an RMSE of 0.0095 on high-order networks, demonstrating superior performance in waveform modeling and transfer function generalization.
@@
— Signalline, transfer function,RC network,generalization,current response, simulation, Transformer, CNN , Deep learning
@@ I Introduction
The dynamic behavior of practical circuits is governed primarily by a small number of “dominant poles.” Typically, researchers select the most significant dozen to several dozen poles—based on energy contribution or frequency‐domain gain thresholds—from an original RC network of thousands or even tens of thousands of orders, and then apply Model Order Reduction (MOR) techniques to reduce the network to the corresponding order [29, 30]. For example, for a 5000th-order RC network, extracting the top 20–50 energy‐contributing poles can reduce the model order to just a few tens of poles, while in most practical cases the frequency response error remains within a few percentage points, thereby greatly lowering model complexity.
Our method seamlessly integrates with the MOR reduction pipeline. On low-order models (e.g., 5th- or 15th-order), the CNN–Transformer architecture achieves a prediction RMSE as low as 0.0098; when applied directly to higher-order reduced models, it still maintains an RMSE of at most 0.0095. Crucially, by exploiting the mathematical fact that any high-order transfer function can be decomposed into a sum of first-order terms, our framework predicts each term’s response independently and then sums them—irrespective of the RC network’s topology—thereby guaranteeing both the universality and consistency of the prediction.
This paper proposes a hybrid architecture combining Transformer and CNN for predicting voltage waveforms in signal lines. Unlike traditional methods such as current source models, driver linear representations, waveform function fitting, or equivalent load capacitance methods, our model does not rely on fixed simplified models of standard-cell drivers or RC loads. Instead, we innovatively introduce the transfer function of the signal line output node as a high-level feature in the modeling process, leveraging the global feature modeling capability of Transformers and the local feature extraction advantages of CNNs to accurately predict sequence responses. By exploiting the decomposability of high-order transfer functions, we construct a generalizable prediction framework that requires training only a small number of sub-models to adapt to high-order RC networks, thus avoiding the cumbersome iterative process and manual modeling involved in traditional methods.
The model adopts a two-stage architecture: the base model predicts the response of the dominant first-order component, while the compensation module further refines the output through a unified correction network. This correction network directly produces point-wise adjustments that are additively applied to the base response, resulting in a complete high-order waveform prediction. Experimental results demonstrate that the proposed method significantly enhances the accuracy of current waveform prediction while effectively reducing modeling complexity and computational overhead.
@@
II Theoretical Background
II-A Decomposition Theorem for High-Order RC Network Transfer Functions
In modern integrated circuit interconnects, the parasitic resistances $R_{i}$ and capacitances $C_{i}$ distributed along metal wires form a linear time-invariant (LTI) RC network whose dynamics govern signal propagation delay, attenuation, and distortion. To characterize this behavior, we use the Laplace-domain transfer function $H(s)$, mapping input voltage or current to output voltage or current. [29][40][41] For an arbitrary LTI RC network, the transfer function has the following essential properties:
II-A1 Network Topology and Mathematical Representation
Assume the network contains $m$ independent energy-storage nodes (capacitors), leading to $m$th-order differential equations. Applying the Laplace transform yields:
$H(s)=\frac{N(s)}{D(s)}=\frac{b_{0}+b_{1}s+\cdots+b_{m-1}s^{m-1}}{a_{0}+a_{1}s+\cdots+a_{m}s^{m}}.$ (1)
The roots of the denominator $D(s)$ are the system poles $p_{i}$, satisfying:
$D(s)=\prod_{i=1}^{m}(s-p_{i}),\quad p_{i}=-\frac{1}{\tau_{i}}\in\mathbb{R}^{-},$ (2)
where $\tau_{i}=R_{i}C_{i}$. The strictly negative real poles follow from the passivity of RC networks.
II-A2 Existence and Uniqueness of Partial Fraction Expansion
By the fundamental theorem of algebra and the Heine–Borel theorem, any strictly proper rational function with distinct poles admits a unique partial fraction expansion[43]:
$H(s)=\sum_{i=1}^{m}\frac{r_{i}}{s-p_{i}},\quad r_{i}=\left.\frac{N(s)}{D^{\prime}(s)}\right|_{s=p_{i}}.$ (3)
This result can be proven rigorously by two methods:
  1. 1.
    Method of Undetermined Coefficients: Multiply both sides by $D(s)$ and equate coefficients of like powers of $s$ to obtain a linear system. The resulting Vandermonde matrix is nonsingular when poles are distinct, guaranteeing a unique solution.
  2. 2.
    Residue Theorem: Integrating $H(s)$ around a contour enclosing all poles and applying Cauchy’s residue theorem yields
[…]
    and analytic continuation guarantees uniqueness.
II-A3 Physical Realizability Conditions
Passivity of RC networks imposes additional constraints[44]:
  • Pole Distribution: All poles lie strictly on the negative real axis, reflecting an overdamped response.
  • Residue Sign: $r_{i}\in\mathbb{R}^{+}$, ensuring positive decaying modes.
  • Impulse Response: $h(t)=\sum_{i=1}^{m}r_{i}e^{p_{i}t}u(t)$, illustrating multi–time-scale coupling of signal propagation.
II-A4 Extension to Repeated Poles
If pole $p_{i}$ has multiplicity $k_{i}>1$ (with $\sum k_{i}=m$), the expansion generalizes to:
$H(s)=\sum_{i=1}^{q}\sum_{j=1}^{k_{i}}\frac{r_{ij}}{(s-p_{i})^{j}},$ (4)
where coefficients are given by higher-order residues:
$r_{ij}=\frac{1}{(k_{i}-j)!}\left.\frac{d^{k_{i}-j}}{ds^{k_{i}-j}}\bigl{[}(s-p_{i})^{k_{i}}H(s)\bigr{]}\right|_{s=p_{i}}.$ (5)
II-A5 Negligible Effect of Repeated Poles
Although repeated poles introduce polynomial factors $t^{j-1}e^{p_{i}t}$ in the time-domain response, in practical on-chip RC networks manufacturing tolerances and layout variations make exact repeated roots unlikely. Even near-coincident poles yield higher-order residues $r_{ij}$ that are typically small, and their $t^{j-1}e^{p_{i}t}$ contributions decay rapidly for $p_{i}<0$. Hence, repeated-pole effects can be safely neglected in most modeling tasks, and single-pole expansions suffice for high-fidelity simulations.[42]
This comprehensive decomposition provides a unified framework for both time-domain and frequency-domain analysis of arbitrary high-order RC interconnect networks.
III Data Acquisition and Preprocessing
III-A Data Source and Simulation Configuration
III-A1 Simulation Environment Setup
@@ III-A2 Stimulation Configuration
III-B Data Preprocessing Methods
III-B1 Data Sources and Feature Definition
@@
  • Transfer Function Features: Instead of directly using RC values from the SPEF netlist, we extract the transfer function from the driver to the output node and perform partial fraction decomposition. The resulting input features are a collection of triplets $(p_{i},k,A_{ik})\in\mathbb{R}\times\mathbb{N}^{+}\times\mathbb{R}$, each representing a real pole $p_{i}$, its multiplicity $k$, and the corresponding residue $A_{ik}$. A fixed number of such terms are selected and sorted by magnitude to form a structured feature vector.
@@ III-B2 Feature Normalization
III-C Transfer Function-Based RC Network Representation
To more compactly and effectively capture the signal behavior of a standard-cell-driven RC interconnect, we propose a transfer function decomposition-based modeling framework. Rather than modeling internal nodes or edges, we directly extract the system-level input-output behavior of the RC network by analyzing its Laplace-domain transfer function.
1. Transfer Function Decomposition
For each RC network, the voltage transfer function from the driver input to the observed output node can be expressed in its partial fraction expansion form:
$H(s)=\sum_{i=1}^{r}\sum_{k=1}^{m_{i}}\frac{A_{ik}}{(s-p_{i})^{k}}$
@@ 1. Transfer Function Decomposition
  • $p_{i}\in\mathbb{R}^{-}$ is the $i$-th real pole (RC networks contain only real negative poles);
  • $m_{i}\in\mathbb{N}^{+}$ is the multiplicity (order) of pole $p_{i}$;
  • $A_{ik}\in\mathbb{R}$ is the residue (numerator coefficient) of the $k$-th order term for pole $p_{i}$;
  • $r$ is the number of distinct poles, and $\sum_{i=1}^{r}m_{i}=n$, the total system order.
This form naturally arises from solving the nodal equations of the RC circuit in the Laplace domain and captures all time constants governing the exponential decay of the transient response.
2. Feature Construction: Pole-Residue Triplets
To enable learning and generalization across circuits of different sizes, we construct a compact, structured feature vector using pole-residue triplets. Each term $\frac{A_{ik}}{(s-p_{i})^{k}}$ is encoded as:
$(p_{i},k,A_{ik})\in\mathbb{R}\times\mathbb{N}^{+}\times\mathbb{R}$
This encoding reflects both the location and order of time constants and the strength of their contribution to the transient behavior. For a fixed representation size, the most dominant terms (by $|A_{ik}|$) are selected, and the rest are zero-padded.
3. Normalization and Sorting Strategy
To ensure numerical stability and consistency across training samples, the following normalization strategy is adopted:
  • All pole locations $p_{i}$ and coefficients $A_{ik}$ are normalized with respect to the largest-magnitude pole $|p_{1}|$;
  • Pole-residue terms are sorted by $|A_{ik}|$ to prioritize dominant dynamics;
  • For circuits with fewer than the maximum number of terms, padding is applied using zeros or a neutral placeholder.
This representation ensures a fixed-length, physically meaningful input encoding suitable for neural network-based learning tasks, preserving the key dynamic modes of the circuit without needing explicit graph structures.
4. Feature Encoding with Repeated Poles
The partial fraction expansion of a transfer function may contain repeated poles. To faithfully encode such components, each term of the form $\frac{A_{ik}}{(s-p_{i})^{k}}$ is represented as a triplet:
$(p_{i},k,A_{ik})\in\mathbb{R}\times\mathbb{N}^{+}\times\mathbb{R}$
@@ 4. Feature Encoding with Repeated Poles
  • $p_{i}$: The real-valued pole location;
  • $k$: The multiplicity (order) of the pole;
  • $A_{ik}$: The numerator coefficient corresponding to the $k$-th order term.
This representation allows the model to explicitly capture the contribution of higher-order modes (e.g., $te^{-pt},t^{2}e^{-pt}$) that arise from repeated poles in the RC transfer function. It also maintains a structured and interpretable format aligned with classical control and circuit theory.
This transfer-function-based modeling scheme transforms the RC network into a system-theoretic representation, providing a powerful and compact alternative to spatially distributed node encodings. It is particularly suitable for waveform prediction tasks where input-output response is governed by the pole-zero distribution.
IV Hybrid Model Architecture Design
IV-A Overall Architecture Overview
@@
IV-B Input Embedding Layer
@@
IV-C CNN Architecture Design
IV-C1 Input Feature Adaptation
@@
IV-C2 Convolutional Feature Extraction
@@
IV-C3 Feature Space Mapping
@@
IV-D Encoder Architecture and Feature Fusion Strategy
@@
IV-D1 Input Feature Preprocessing
@@
IV-D2 Multi-Head Self-Attention Mechanism
@@
IV-D3 Residual Connection and Layer Normalization
@@
IV-D4 Feedforward Neural Network
@@
IV-D5 Second Residual Connection and Normalization
@@ IV-D6 Encoder Output
IV-D7 Feature Fusion via Dot Product
@@
IV-D8 Decoder Input Projection
@@
IV-E Transformer Decoder Architecture
@@ IV-E1 Input Definition
IV-E2 Masked Multi-Head Self-Attention
@@
IV-E3 Encoder-Decoder Cross-Attention
@@
IV-E4 Feedforward Neural Network
@@ IV-E5 Summary
V Results
V-A Scalability Across Network Sizes
@@
V-B Accuracy of Voltage Response Prediction
@@
V-C Training time and Infer time Comparison
@@
VI Future Challenges and Directions for our model
VII Future Challenges and Directions
@@
VII-A Limited Generalization Across Topologies
@@
VII-B Quadratic Complexity of Transformer Inference
@@
VII-C Temporal Alignment and Stability Under Noise
@@
VII-D Lack of Physical Interpretability and Control